OpenStudy (anonymous):

There is an equal number of boys and girls in the 9th grade class. The ratio of the average heights of boys to girls is 12 to 11. If the average height of the girls in the class is 66 inches, what is the average height, in inches, of the boys in the class?

OpenStudy (anonymous):

b:f=12=11 11x=66 x=6 x(12)=72

OpenStudy (charlotte123):

12:11 ^ * 6 ----> 66 12 * 6 = 72

OpenStudy (charlotte123):

72 Inches <----

OpenStudy (calculusxy):

Solve a proportions: \[\frac{ boys }{ girls }=\frac{ 12 }{ 11 }=\frac{ x }{ 66 }\]

OpenStudy (charlotte123):

^

OpenStudy (calculusxy):

Now you can cross multiply to get the value of x.